Solve the differential equation by variation of parameters 1 x2y + x y- y=; х
Answer #1

criven diff. equ x²y + xy-y = 14 first we have to find C.F. [complemantry Enction x²y try-y=0 Let x=e2 - 2 = loga Then eq23 vin 2 x2 Again diff. @ w. h. t. n y = A + 20 - B using @ 2[A.* - ] + x[A-]-[Ax+]=+ A, R² of .B, a AK-B-An- - Z - A % = 1/